OpenStudy (anonymous):

Find the equation of a line going through points (-3, 4) and (0, 5). Write it in slope-intercept form. Show your work. Please help me!!!!!

OpenStudy (amistre64):

subtract one point from the other to start with

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

did you mean -3 and 4 to subtract?

OpenStudy (amistre64):

good question. no, what i like to do is stack the points and subtract like parts like this (-3, 4) -( 0, 5) -------- -3, -1

OpenStudy (amistre64):

the result gives me what I need to make a slope with: slope = y/x or -1/-3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is the formula for this y=mx+b

OpenStudy (amistre64):

the end result will be yes

OpenStudy (amistre64):

m is where we put our slope that we just determined; m=1/3

OpenStudy (amistre64):

b can be found a couple of ways; the most straight forward is to use one of your points to find it y = 1/3 x + b ; spose we use (0,5) 5 = 1/3 * 0 + b 5 = 0 + b

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

thats it really

OpenStudy (amistre64):

you know m and b ... so we fill them in to make the desired equation

OpenStudy (anonymous):

so what i should do is plug in y= mx+b which would be Y=1/3(0)+b

OpenStudy (amistre64):

y = mx + b m = 1/3 and b = 5 y = 1/3 x + 5

OpenStudy (anonymous):

now if i plugged in the (-3,4) it would stil work out the same?

OpenStudy (amistre64):

yes if you used the other point it would have worked out the same since both points are on the same line; and the only way that can happen is if the slope of the line is created from the 2 points

OpenStudy (amistre64):

y = 1/3 x + b 4 = 1/3 * -3 + b 4 = -1 + b 5 = b

OpenStudy (amistre64):

y = 1/3 x + 5

OpenStudy (anonymous):

cool thanks for helping me understand it better.

OpenStudy (amistre64):

good luck :)